OpenStudy (anonymous):

Two pipes fill a storage tank in 9 hours. The larger pipe fills the tank three times as fast as the smaller one. How long would it take the larger pipe to fill the tank alone

OpenStudy (anonymous):

let the smaller pipe rate be v so the larger will be 3v (v+3v) * 9 = 1 36v = 1 v=1/36 so in order to find t the time that it would take for the larger pipe all we haveto do is 3v * t = 1 you plug here the v that we found up there

OpenStudy (anonymous):

21 hours?

OpenStudy (anonymous):

or 18. @Coolsector

OpenStudy (anonymous):

lets do this together : 3v*t=1 3 * (1/36) * t = 1 (1/12) * t = 1 t = 12